OpenStudy (anonymous):

solve the equation using the square root property. I need to understand how to break it down to get the answer. -5(x-6)2-15=-30 the 2 is square

OpenStudy (phi):

use ^ to show a power: -5(x-6)^2-15=-30

OpenStudy (phi):

do you know you could add 15 to both sides of the equation, right? \[ -5(x-6)^2-15+15=-30+15\] can you simplify this?

OpenStudy (anonymous):

then divide 5 to both sides?

OpenStudy (phi):

divide by -5

OpenStudy (anonymous):

now I have (x-6) = 3

OpenStudy (anonymous):

add 6 to each side

OpenStudy (phi):

what happened to the ^2 ?

OpenStudy (phi):

it should be (x-6)^2 = 3

OpenStudy (phi):

now you can do the square root on both sides

OpenStudy (anonymous):

x=9

OpenStudy (phi):

no. -5(x-6)^2-15=-30 becomes (x-6)^2 = 3 you must take the square root of both sides

OpenStudy (anonymous):

x=3 or x= -3

OpenStudy (anonymous):

are you still there?

OpenStudy (phi):

do you know how to take the square root of (x-6)^2 ?

OpenStudy (anonymous):

I thought I did, but can you show me please?

OpenStudy (anonymous):

is it x=2 or x-2

OpenStudy (phi):

\[ \sqrt{ x^2} = x\] \[ \sqrt{ (x+2)^2} = x+2\]